A manufacture of rayon wants to compare that the yield strength of 5 11⋅ kg/ 2 mm is 

met or not at 5% level of significance. The manufacturer draws a sample and 

calculates the mean to be 8 12⋅ kg/ 2 mm and the standard derivation is known to be 

2⋅0kg/ mm .

2

 Carry out the statistical test appropriate for this.


1
Expert's answer
2021-05-19T07:57:46-0400

Null and Alternative Hypotheses

The following null and alternative hypotheses need to be tested:

"H_0: \mu=5.11\ kg/mm^2"

"H_1: \mu\not=5.11\ kg/mm^2"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha=0.05" and "df=n-1=30-1=29" degrees of freedom.

The critical value for a two-tailed test is "t_c=2.04523."  

The rejection region for this two-tailed test is "R=\{t:|t|>2.04523\}."


The "t" - statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{8.12-5.11}{2.0/\sqrt{30}}\approx8.24322"

Since it is observed that "|t|=8.24322>2.04523=|t_c|," it is then concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"  is different than "5.11," at the "\alpha=0.05" significance level.


Using the P-value approach: The p-value for two-tailed, the significance level  "\alpha=0.05, t=8.24322" and "df=29" degrees of freedom is "p<0.00001,"

and since "p<0.00001<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"  is different than "5.11," at the "\alpha=0.05" significance level.


Therefore, there is enough evidence to claim that the yield strength is different than "5.11," at the "\alpha=0.05" significance level.